# Toward Signed Distance Function based Metamaterial Design: Neural Operator Transformer for Forward Prediction and Diffusion Model for Inverse Design

### v2

@@ Abstract
Keywords: Diffusion model; Inverse design; Neural operator transformer; Arbitrary geometry; Metamaterial
@@ 2.1 Data generation
The generated geometries are then imported into the commercial finite element code ABAQUS® 2024 to compute the macroscopic stress-strain curve and the local solution fields. We fix the bottom of the unit cell along the y-direction ($u_{y}=0$) and deform the unit cell by applying a vertical displacement to the top edge of the unit cell with a maximum applied compressive strain of 0.2. Periodic boundary conditions are applied along the left and right sides of the unit cell. The horizontal displacement of the node nearest to the center of the unit cell is set at $u_{x}=0$ to remove the rigid body motion. Self-contact is enabled to prevent the unit cell from collapsing during compression. Simulation details are provided in the methods Section 4.2.
Based on this method, we randomly generate 10k geometries and their corresponding stress-strain curves and solution fields (e.g., Mises stress, displacement). Extended Data Fig. 1 illustrates the generated dataset, where (a) shows the volume fraction of the 10k samples, (b) displays 64 examples of the generated geometries, and (c) presents their corresponding stress-strain curves.
@@ 2.2 Forward prediction
The performance of the forward model for predicting the stress-strain curve from arbitrary unit cell geometries is illustrated in Fig. 3, which shows a comparison between the true and predicted stress-strain curves of the test data, arranged in terms of $L_{2}$ relative error from best (left) to the 99th percentile (right). Further performance details for predicting the stress-strain curve are presented in Fig. S4, which shows the MSE loss of training history and the $L_{2}$ relative error distribution of the test data. The overall mean $L_{2}$ relative error of the test data is 2.6% with a standard deviation of 2.4%, indicating the model’s high accuracy in predicting the stress-strain curve.
We also demonstrate in Fig. 4 the performance of the forward model in predicting the Mises stress and displacement fields for the median case of the test data. The first row displays the true (FE-predicted) Mises stress field plotted on the true deformed shape at different strain steps. The second row shows the predicted (by the NOT) Mises stress under the predicted deformed shape, while the third row presents the corresponding absolute error in Mises stress plotted on the true deformed shape of the unit cell. The $L_{2}$ relative error for the median case is 10.1%. Additionally, we present the best and worst cases of the test data in Fig. S6 and Fig. S7, respectively. The MSE loss of the training history and the $L_{2}$ relative error distribution of the test data for Mises stress and displacement solutions are shown in Fig. S5. The overall mean $L_{2}$ relative error of the test data is 10.3% with a standard deviation of 4.6%, indicating the model’s high accuracy in predicting the solution fields for arbitrary geometries.
@@ 2.3 Inverse design by diffusion model
In this section, we describe the classifier-free guided diffusion model [40] trained to design the micro-structure of the unit cell of the metamaterial that can achieve a target stress-strain curve. The label data $y$ is the stress-strain curve obtained from the FE simulation of the corresponding geometry, represented by SDF ($x_{0}$).
@@
Such a generative diffusion model can perform one-to-many designs, where giving a single target stress-strain curve generates multiple unit cell geometries that satisfy the target stress-strain curve. We demonstrate the trained diffusion model for such one-to-many material architecture designs by randomly selecting two target stress-strain curves from the test dataset and feeding them into the inverse diffusion model for 200 geometry solutions. The generated geometries are then fed into the forward model to predict the stress-strain curve. We select the four best design results in terms of $L_{2}$ relative error between the target and prediction of the forward model, as shown in Extended Data Fig. 2(c-f). The comparison between the target and prediction of the forward model is shown in Extended Data Fig. 2(a), with the $L_{2}$ relative error in the legend bracket. The corresponding stress-strain curves obtained using Abaqus simulation are shown in Extended Data Fig. 2(b). The results show that the designed geometries can achieve the target stress-strain curve with high accuracy. With the designed geometry, we can further predict the solution fields (e.g., Mises stress, displacement) at different strain steps using our second forward NOT. The comparison of solution fields for the generated geometry of Extended Data Fig. 2(A)(f) between the true and predicted results is shown in Extended Data Fig. 3.
@@
Figure 6: Comparison of the Mises stress and displacement field predictions with FE ground truth at different strain $\varepsilon$ steps for the designed structure in Fig. 5(A)(e). The first row shows the true Mises stress on the true deformed shape, the second row shows the predicted Mises stress on the predicted deformed shape, and the third row shows the absolute error in Mises stress on the true deformed shape.
While our developed diffusion model performs well on the test dataset for inverse retrieval, real-world applications often demand on-the-fly inverse design of custom-defined material properties with high fidelity. Therefore, instead of using stress-strain curves from FE simulations as shown in Extended Data Fig. 2, a robust inverse design model should be able to generate possible designs from custom and predefined stress-strain curves. Thus, we now use the Ramberg-Osgood equation [41] to define the target stress-strain curve. The Ramberg-Osgood equation is defined as
@@
Using the Ramberg-Osgood equations, we generate two target stress-strain curves with different material properties, as shown in Fig. 5(A) and (B). Both targets use the same hardening exponent $n=10$ and coefficient $\alpha=0.002$, but differ in Young’s modulus and reference yield stress. Case (A) is obtained using Eq. 7 with a Young’s modulus $E=800$ MPa and yield stress $\sigma_{0}=30$ MPa, while Case (B) is derived from Eq. 8 with $E=1000$ MPa, $\sigma_{0}=40$ MPa, $E^{\prime}=500$ MPa, and $\varepsilon_{c}=0.156$. These target curves are fed into the inverse diffusion model to generate 500 geometry solutions, which are then evaluated using the forward model to predict the stress-strain curves. The four best designs based on the $L_{2}$ relative error are shown in Fig. 5 (c-f). Fig. 5(a) and (b) compare the target and designed stress-strain curves for the four geometries using the forward model and FE simulations, respectively, with the $L_{2}$ relative error indicated in the curve legend. As shown in Fig. 5(a), the designed geometries closely match the custom-defined stress-strain curves. The FE simulations of the designed geometries, shown in Fig. 5(b), confirm the accuracy of the designs. Additionally, we predict the solution fields for the designed geometry in Fig. 5(A)(e) at different strain steps, with results shown in Fig. 6. For case (B), the solution fields for the designed geometry (e) are shown in Extended Data Fig. 5. These results demonstrate that the designed geometries achieve the target stress-strain curves and accurately predict the solution fields.
3 Discussion
@@ 4.4 Training protocol
Before training, we normalize the SDF, stress, and displacement to zero mean and unit variance. The dataset is split into 80% for training and 20% for testing. The mean square error is used as the loss function. We implement this work using the PyTorch framework, and training is performed on a single NVIDIA H100 GPU with 90GB memory on the DeltaAI machine available at NCSA. We use the Adam optimizer [44] with an initial learning rate of $0.001$, leveraging its adaptive moment estimation for stable and efficient convergence. To further enhance training stability and adapt the learning rate dynamically, we incorporate the ReduceLROnPlateau scheduler, which reduces the learning rate when the validation loss stagnates. For padding query points, a mask mechanism is applied to exclude them from the loss calculation. The loss value evolution during training is shown in Fig. S4 and Fig. S5. Detailed training information is provided in the Supplementary Information. Specifically, the computational efficiency is discussed in Supplementary Section S5.
@@ 7 Acknowledgements
The authors would like to thank the National Center for Supercomputing Applications (NCSA) at the University of Illinois Urbana-Champaign, and particularly its Research Computing Directorate, Industry Program, and Center for Artificial Intelligence Innovation (CAII) for support and hardware resources. This research is a part of the Delta research computing project, which is supported by the National Science Foundation (award OCI 2005572) and the State of Illinois, as well as the Illinois Computes program supported by the University of Illinois Urbana-Champaign and the University of Illinois System.

### v3

@@ Abstract
Keywords: Diffusion model; Inverse design; Neural operator transformer; Arbitrary geometry; Metamaterial; Generative AI
@@ 2.1 Data generation
The generated geometries are then imported into the commercial finite element code ABAQUS® 2024 to compute the macroscopic stress-strain curve and the local solution fields. We fix the bottom of the unit cell along the y-direction ($u_{y}=0$) and deform the unit cell by applying a vertical displacement to the top edge of the unit cell with a maximum applied compressive strain of 0.2. Periodic boundary conditions are applied along the left and right sides of the unit cell. Self-contact is enabled to prevent the unit cell from collapsing during compression. Simulation details are provided in the methods Section 3.2.
Figure 2: Illustration of training samples. (a) Volume fraction distribution of the dataset. (b) 64 typical microstructure geometries. (c) Corresponding stress-strain curves for the geometries shown in (b).
Based on this method, we randomly generate 10k geometries and their corresponding stress-strain curves and solution fields (e.g., Mises stress, displacement). Fig. 2 illustrates the generated dataset, where (a) shows the volume fraction of the 10k samples, (b) displays 64 examples of the generated geometries, and (c) presents their corresponding stress-strain curves.
@@ 2.2 Forward prediction
The performance of the forward model for predicting the stress-strain curve from arbitrary unit cell geometries is illustrated in Fig. 4, which shows a comparison between the true and predicted stress-strain curves of the test data, arranged in terms of $L_{2}$ relative error from best (left) to the 99th percentile (right). Further performance details for predicting the stress-strain curve are presented in Supplementary Information (Fig. S4), which shows the MSE loss of training history and the $L_{2}$ relative error distribution of the test data. The overall mean $L_{2}$ relative error of the test data is 2.6% with a standard deviation of 2.4%, indicating the model’s high accuracy in predicting the stress-strain curve.
We also demonstrate in Fig. 5 the performance of the forward model in predicting the Mises stress and displacement fields for the median case of the test data. The first row displays the true (FE-predicted) Mises stress field plotted on the true deformed shape at different strain steps. The second row shows the predicted (by the NOT) Mises stress under the predicted deformed shape, while the third row presents the corresponding absolute error in Mises stress plotted on the true deformed shape of the unit cell. The $L_{2}$ relative error for the median case is 10.1%. Additionally, we present the best and worst cases of the test data in Supplementary Information (Fig. S6 and Fig. S7). The MSE loss of the training history and the $L_{2}$ relative error distribution of the test data for Mises stress and displacement solutions are shown in Supplementary Information Fig. S5. The overall mean $L_{2}$ relative error of the test data is 10.3% with a standard deviation of 4.6%, indicating the model’s high accuracy in predicting the solution fields for arbitrary geometries.
@@ 2.3 Inverse design by diffusion model
In this section, we first describe the classifier-free guided diffusion model [40] trained to design the micro-structure of the unit cell of the metamaterial that can achieve a target stress-strain curve. The label data $y$ is the stress-strain curve obtained from the FE simulation of the corresponding geometry, represented by SDF ($x_{0}$).
@@
Figure 6: One-to-many unit cell design. Two targets, (A) and (B), of stress-strain curves are randomly selected from the test dataset and fed into the inverse diffusion model with guidance weight $w=10$ to generate 200 geometry solutions. These solutions are then evaluated using the forward model to predict the stress-strain curves. We select the 4 best design results based on the $L_{2}$ relative error between the target and the forward model predictions, shown in (c-f). (a) shows the comparison between the target and the predictions of (c-f) using the forward model, with the $L_{2}$ relative error indicated in the legend. (b) shows the comparison between the target and the stress-strain curves obtained using Abaqus simulation for (c-f).
Figure 7: Comparison between the Mises stress and displacement field predictions with the corresponding FE ground truth at different strain $\varepsilon$ steps for the designed structure of Fig. 6(Af). The first row shows the true Mises stress field displayed on the true deformed shape; the second row shows the predicted Mises stress field onto the predicted deformed shape; and the third row shows the absolute error in Mises stress plotted on the true deformed shape.
Figure 8: Comparison between the Mises stress and displacement field predictions with the corresponding FE ground truth at different strain $\varepsilon$ steps for the designed structure of Fig. 6(Bd). The first row shows the true Mises stress under the true deformed shape; the second row shows the predicted Mises stress under the predicted deformed shape; and the third row shows the absolute error of the Mises stress under the true deformed shape.
Such a generative diffusion model can perform one-to-many designs, where giving a single target stress-strain curve generates multiple unit cell geometries that satisfy the target stress-strain curve. We demonstrate the trained diffusion model for such one-to-many material architecture designs by randomly selecting two target stress-strain curves from the test dataset and feeding them into the inverse diffusion model for 200 geometry solutions. The generated geometries are then fed into the forward model to predict the stress-strain curve. We select the four best design results in terms of $L_{2}$ relative error between the target and prediction of the forward model, as shown in Fig. 6(c-f). The comparison between the target and prediction of the forward model is shown in Fig. 6(a), with the $L_{2}$ relative error in the legend bracket. The corresponding stress-strain curves obtained using Abaqus simulation are shown in Fig. 6(b). The results show that the designed geometries can achieve the target stress-strain curve with high accuracy. With the designed geometry, we can further predict the solution fields (e.g., Mises stress, displacement) at different strain steps using our second forward NOT. The comparison of solution fields between the true and predicted results for the generated geometry of Fig. 6(Af) and Fig. 6(Bd) are shown in Fig. 7 and Fig. 8, respectively.
@@
Figure 10: Comparison of the Mises stress and displacement field predictions with FE ground truth at different strain $\varepsilon$ steps for the designed structure in Fig. 9(Ae). The first row shows the true Mises stress on the true deformed shape, the second row shows the predicted Mises stress on the predicted deformed shape, and the third row shows the absolute error in Mises stress on the true deformed shape.
Figure 11: Comparison between the Mises stress and displacement field predictions with the corresponding FE ground truth at different strain $\varepsilon$ steps for the designed structure in Fig. 9(Be). The first row shows the true Mises stress under the true deformed shape; the second row shows the predicted Mises stress under the predicted deformed shape; and the third row shows the absolute error of the Mises stress under the true deformed shape.
While our developed diffusion model performs well on the test dataset for inverse retrieval, real-world applications often demand on-the-fly inverse design of custom-defined material properties with high fidelity. Therefore, instead of using stress-strain curves from FE simulations as shown in Fig. 6, a robust inverse design model should be able to generate possible designs from custom and predefined stress-strain curves. Thus, we now use the Ramberg-Osgood equation [41] to define the target stress-strain curve. The Ramberg-Osgood equation is defined as
@@
Using the Ramberg-Osgood equations, we generate two target stress-strain curves with different material properties, as shown in Fig. 9(A) and (B). Both targets use the same hardening exponent $n=10$ and coefficient $\alpha=0.002$, but differ in Young’s modulus and reference yield stress. Case (A) is obtained using Eq. 7 with a Young’s modulus $E=800$ MPa and yield stress $\sigma_{0}=30$ MPa, while Case (B) is derived from Eq. 8 with $E=1000$ MPa, $\sigma_{0}=40$ MPa, $E^{\prime}=500$ MPa, and $\varepsilon_{c}=0.156$. These target curves are fed into the inverse diffusion model to generate 500 geometry solutions, which are then evaluated using the forward model to predict the stress-strain curves. The four best designs based on the $L_{2}$ relative error are shown in Fig. 9 (c-f). Fig. 9(a) and (b) compare the target and designed stress-strain curves for the four geometries using the forward model and FE simulations, respectively, with the $L_{2}$ relative error indicated in the curve legend. As shown in Fig. 9(a), the designed geometries closely match the custom-defined stress-strain curves. The FE simulations of the designed geometries, shown in Fig. 9(b), confirm the accuracy of the designs. Additionally, we predict the solution fields for the designed geometries Fig. 9(Ae) and Fig. 9(Be) at different strain steps, with results shown in Fig. 10 and Fig. 11, respectively. These results demonstrate that the designed geometries achieve the target stress-strain curves and accurately predict the solution fields.
@@ 3.4 Training protocol
Before training, we normalize the SDF, stress, and displacement to zero mean and unit variance. The dataset is split into 80% for training and 20% for testing. The mean square error is used as the loss function. We implement this work using the PyTorch framework, and training is performed on a single NVIDIA H100 GPU with 90GB memory on the DeltaAI machine available at NCSA. We use the Adam optimizer [44] with an initial learning rate of $0.001$, leveraging its adaptive moment estimation for stable and efficient convergence. To further enhance training stability and adapt the learning rate dynamically, we incorporate the ReduceLROnPlateau scheduler, which reduces the learning rate when the validation loss stagnates. For padding query points, a mask mechanism is applied to exclude them from the loss calculation. The loss value evolution during training is shown in Supplementary Information (Fig. S4 and Fig. S5.) Detailed training information is provided in the Supplementary Information. Specifically, the computational efficiency is discussed in Supplementary Section S5.
@@ 7 Acknowledgements
The authors would like to thank the National Center for Supercomputing Applications (NCSA) at the University of Illinois Urbana-Champaign, and particularly its Research Computing Directorate, Industry Program, and Center for Artificial Intelligence Innovation (CAII) for support and hardware resources. This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.